OpenStudy (anonymous):

On the graph of the equation 3x – 2y = –18, what is the value of the y-intercept?

OpenStudy (anonymous):

\[3x-2y=-18\]\[-2y=-3x-18\] \[y=\frac{-3x}{-2} -\frac{18}{-2} \] the y intercept is \[ b=9\]

OpenStudy (anonymous):

3x – 2y = –18 y-intercept -> x is 0 -2y=-18 y=9 (0,9)